OpenStudy (anonymous):

Sam is riding his bicycle. He rides at a speed of 11 kilometers per hour for 17.6 kilometers. For how many hours does he ride?

OpenStudy (anonymous):

Use proportions to solve this problem. 11 km per hours means 11km divided by one hour. So write it like this: \[\frac{ 11km }{ 1 hour }\] Now write the other side of the proportion. Since kilometer is in the numerator and hours are in the denominator, the other side of the proportion must have the these units in the same position. So our proportion will look like this: \[\frac{ 11km }{ 1 hour }=\frac{ 17.6km }{ ?hours }\] Solve the proportion by cross multiplying and then dividing by the other number. So we have this: \[(1hour*17.6km)\div11km=1.6 hours\]